Investigating a complaint from a buyer that there is short-weight selling, a manufacturer takes a random sample of twenty-five 32 g cans of coffee from a large shipment and finds that the mean weight is 31 g with a standard deviation of 0.6 g. Is there evidence of short-weighing at the 0.01 level significance?


1
Expert's answer
2022-05-11T17:51:10-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge32"

"H_a:\mu<32"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=34" degrees of freedom, and the critical value for a left-tailed test is "t_c = -2.44115."

The rejection region for this left-tailed test is "R = \{t: t<-2.44115\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{31-32}{0.6/\sqrt{35}}\approx-9.8601"


Since it is observed that "t=-9.8601< -2.44115=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed, "df=34" degrees of freedom, "t=-9.8601" is "p=0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than 32, at the "\alpha = 0.01" significance level.